Problem 14-10 Taxes and WACC [LO3]

Lannister Manufacturing has a target debt-equity ratio of .45. Its cost of equity is 11 percent, and its cost of debt is 6 percent. If the tax rate is 25 percent, what is the company’s WACC? (

Debt:

8,000 5.7 percent coupon bonds outstanding, $1,000 par value, 23 years to maturity, selling for 105 percent of par; the bonds make semiannual payments.

  Common stock: 410,000 shares outstanding, selling for $59 per share; the beta is 1.10.
  Preferred stock:

17,500 shares of 3.5 percent preferred stock outstanding, currently selling for $80 per share. The par value is $100 per share.

  Market: 7 percent market risk premium and 4.5 percent risk-free rate.

What is the company's WACC?

WACC %
Answer #1

WACC of Lannister Corporation =Weight of Equity*Cost of Equity+Weight of Debt*Cost of Debt*(1-Tax rate)
=1/(1+0.45)*11%+0.45/(1+0.45)*6%*(1-25%) =8.98%

Value of debt =Number of Bonds*105%*Par Value =8000*105%*1000 =8,400,000
Semi annual coupon =5.7%*1000/2 =28.50
Number of Periods =23*2 =46
Par Value =1000
Price =1000*105%=1050
YTM using excel formula =2*RATE(46,28.50,-1050,1000) =5.3206%

Value of Equity =Number of Shares *Price of Share =410000*59 =24,190,000
Cost of equity =Risk free Rate+Beta*(Market Return-Risk Free Rate) =4.5%+1.10*7% =12.20%

Value of Preferred Stock =17500*80 =1,400,000
Cost of Preferred Stock =Dividend/Price =3.5%*100/80 =4.3750%

Total Value =8,400,000+24,190,000+1,400,000 =33,990,000

WACC =Weight of Equity*Cost of Equity+Weight of Preferred Stock*Cost of Preferred Stock+Weight of Debt*Cost of Debt*(1-Tax Rate) =24190000/33990000*12.20%+1400000/33990000*4.3750%+8400000/33990000*5.3206%*(1-25%) =9.85%